OpenStudy (anonymous):

solve the equation 6b+4=2(2b+1)

OpenStudy (ajprincess):

vat do u get?

OpenStudy (anonymous):

b=1.5

OpenStudy (ajprincess):

no. try again.

OpenStudy (anonymous):

okay

OpenStudy (ajprincess):

did u try? vat r u gettng nw?

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

b=0.6

OpenStudy (ajprincess):

6b+4=2(2b+1) 6b+4=2*2b+2*1 6b+4=4b+2 Nw try.

OpenStudy (anonymous):

10b=6.....?

OpenStudy (ajprincess):

6b+4=4b+2 6b+4-4=4b+2-4 6b=4b-2 6b-4b=4b-2-4b 2b=-2

OpenStudy (anonymous):

b=-1

OpenStudy (ajprincess):

ya right.

OpenStudy (anonymous):

thank you

OpenStudy (ajprincess):

u r welcome